# SafeGen: Mitigating Unsafe Content Generation in Text-to-Image Models

### v1

@@ 4.4 System Integration
From a systematic view, the self-attention layer regulation method of SafeGen can seamlessly integrate other defenses as a complement. Our design boosts the compliance of unconditionally vision-only (i.e., text-agnostic) process within the classifier-free guidance (as illustrated in Equation (2)) without interfering with the conditionally text-dependent process. Hence, our method can collaborate with internal text-dependent countermeasures, particularly the guidance-based SLD [13] to provide stronger protection that ensures safety for both conditional $\epsilon_{\mathtt{U^{*}}(z_{t},c,t)}$ and unconditional $\epsilon_{\mathtt{U^{*}}(z_{t},t)}$ denoising diffusion processes. Similarly, our method aligns well with ESD [14]. Our evaluation of the complementary perspective is elaborated in §6.3.
@@
Figure 9: SafeGen effectively mitigates sexually explicit content yet retains the high-fidelity benign creation.
5 Experiment Setup
@@
We implement SafeGen using Python 3.8 and Pytorch 1.12 on a Ubuntu 22.04 server. All experiments are performed using an A100-40GB GPU (NVIDIA). SafeGen merely edits the self-attention layers of the U-Net module in Stable Diffusion models and can integrate with text-dependent methods. We follow previous work [16, 14] to use Stable Diffusion (version 1.4) unless specified, more details including hyperparameters can be found in Appendix C.
5.1 Baselines
@@
We compare SafeGen with eight baselines, each exemplifying the latest anti-NSFW countermeasures. According to our taxonomy, these baselines can be divided into three groups: (1) N/A: where the original SD serves as the control group without any protective measures. (2) External Mitigation: involving safety filters to block inadvertently generated NSFW images [11], although susceptible to bypassing by adversarial prompts; alternatively, conducting the training data censorship to minimize exposure to NSFW content and retraining the SD model using the censored data [12], requiring substantial computation resources. (3) Internal Mitigation: involving representative text-dependent methods that steer the denoising diffusion process away from NSFW areas. Existing work either adopts guidance-based [13] or model weights modification-based [14], but both are text-dependent and need predefined NSFW concepts. The details of these baselines are listed as follows:
@@
  • [N/A] SD: Stable Diffusion [3], we follow previous work [16, 14] to use the officially provided Stable Diffusion V1.4 [3].
@@
  • [External Filter] SD with safety filter: we use the officially released image-based safety checker [11] to examine its performance in detecting unsafe images.
@@
  • [External Censorship] SD-V2.1: Stable Diffusion V2.1, we use the official version [12], which is retrained on a large-scale dataset censored by external filters.
@@
  • [Internal Text-Dependent] SLD: Safe Latent Diffusion, we adopt the officially pre-trained model [43]; our configuration examines its four safety levels, i.e., weak, medium, strong, and max.
@@
  • [Internal Text-Dependent] ESD: Erased Stable Diffusion, we follow its instruction [14], which erases the concept “nudity” and trains the model for 1000 epochs with learning rate 1e-5.
5.2 Evaluating Metrics
@@
We evaluate a T2I model’s ability in safe generation from two perspectives: (1) NSFW content removal, which is used to evaluate the model’s effectiveness in reducing NSFW content; and (2) benign content preservation, which is used to evaluate the model’s ability to preserve the high quality benign content generation. We use the following four metrics.
@@
  • [NSFW Removal] NRR: The NSFW removal rate (NRR) is a metric for calculating the effectiveness of a T2I model in moderating NSFW content from images compared with the [N/A] original SD-V1.4 without safety mechanisms. NRR is calculated by NudeNet [44]. For each generated image, NudeNet [44] first identifies exposed body parts like breasts or bellies, it then aggregates the number of all identified parts as the total number of NSFW parts found in the image. The NRR refers to the difference in the number of detected NSFW parts between SafeGen or baseline methods and the SD-V1.4 model, a higher NRR indicates more effectiveness, meaning that more identified NSFW parts generated by the SD-V1.4 model have been successfully moderated. We first illustrate the overall effectiveness of SafeGen on different datasets by showing the NRR on total identified parts, we then show that SafeGen continuously outperforms baselines with a higher NRR on different NSFW parts.
@@
  • [NSFW Removal/Benign Preservation] CLIP Score: CLIP enables machines to interpret the relationships between images and their associated captions. Based on its significant zero-shot transferability, for each prompt, CLIP score computes the average cosine similarity between the given CLIP text embedding and its generated CLIP image embedding. In terms of benign generation, a higher score denotes that the T2I model can faithfully reflect the user’s prompt by way of images. In contrast, when confronted with a sexually explicit prompt, a lower score indicates the tested T2I model is safer as its generation deviates from the adversary’s desire.
@@
  • [Benign Preservation] LPIPS Score: The Learned Perceptual Image Patch Similarity (LPIPS) score [45] is another metric for evaluating the fidelity of generated images. LPIPS works by mimicking human visual perception, it captures the difference between detailed image features, such as texture and color. A lower score on the LPIPS score indicates that the two images are more visually similar.
@@
  • [Benign Preservation] FID Score: Different from the LPIPS focuses on the detailed comparison between two images, the Frechet Inception Distance (FID) score [46] is a metric to compare the quality and fidelity between a set of created images and the other set of reference images. We evaluate the benign generated images’ quality of T2I models based on FID scores. A lower score on the FID score means that the two image sets’ distributions are more similar.
5.3 Adversarial and Benign Prompt Benchmark
@@
Our methodology is evaluated using a comprehensive benchmark that encompasses four different prompt datasets. To assess the effectiveness of SafeGen in reducing NSFW generation, we utilize three adversarial prompt datasets, including the widely tested I2P dataset, along with our constructed SneakyPrompt and NSFW-56k datasets. Additionally, we employ a benign prompt dataset, COCO-2017, to evaluate SafeGen’s ability in maintaining high-fidelity benign generation.
@@
  • I2P: Inappropriate Image Prompts [18] consist of manually-tailored NSFW text prompts on lexica.art, from which we select all sex-related prompts, resulting in a total of 931 samples.
@@
  • SneakyPrompt: To evaluate the effectiveness of SafeGen against adaptive adversaries capable of generating sexually connotated prompts via optimization, we reproduce SneakyPrompt [16] and provide two versions of re-use prompt: i.e., SneakyPrompt-N with natural words, and SneakyPrompt-P with pseudo words.
@@
  • NSFW-56k: This dataset encompasses the real-world instances of sexual exposure, as it is constructed by the crawling pornographic image datasets [47]. We follow the CLIP Interrogator [48] to construct a large image-to-prompt dataset. For each image, we first use BLIP2 [19] to get multiple candidate text captions, then choose the best one with the highest CLIP score [29] between image and text captions. This dataset consists of 56k prompts.
@@
  • COCO-25k: We use the validation subset of MS COCO-2017 dataset. Each image within this dataset has been correspondingly captioned by five human annotators, and the associated images were utilized as reference to gauge image fidelity.
6 Evaluation
@@
  • [RQ1] How effective is SafeGen in mitigating the NSFW generation from different types of adversarial prompts?
@@
6.1 RQ1: NSFW Generation Mitigation
We compare SafeGen with eight baselines, i.e., SD with different strategies of anti-NSFW countermeasures, and show SafeGen outperforms all baselines in mitigating NSFW generation across two key metrics. First, we use the NSFW removing rate (NRR) to show that SafeGen is effective in removing the NSFW content, e.g., explicit body parts, among different adversarial prompts. Second, we use the CLIP score to show that SafeGen can reduce the text-to-image alignment between various adversarial prompts and their generation.
@@ 6.1 RQ1: NSFW Generation Mitigation
Mitigation Method NRR (NSFW Removal Rate) normal-$\uparrow$
@@
N/A Original SD - - - -
External Censorship & Filter SD-V2.1 58.8% 45.9% 43.8% 57.0%
Safety Filter 67.9% 67.1% 70.8% 70.0%
Internal Text-dependent ESD 80.3% 80.6% 59.7% 67.6%
SLD (Max) 72.7% 69.3% 75.2% 60.7%
SLD (Strong) 53.3% 44.2% 63.6% 41.5%
SLD (Medium) 27.6% 22.2% 39.4% 20.6%
SLD (Weak) 10.9% 7.5% 9.8% 5.6%
Text-agnostic SafeGen (Ours) 97.8% 97.6% 91.6% 99.1%
Table I: [RQ1-NRR] Performance of SafeGen on NSFW removal rate compared with baselines on different adversarial prompt datasets.
@@
Figure 10: [RQ1-NRR] We show the NSFW removal rate (NRR) in the generated images classified as NSFW by NudeNet [44] compared to that from the original SD-V1.4 model. Our approach effectively reduces the NSFW content and outperforms all prior methods, i.e., SD-V2.1 [12], ESD [14], SLD [13] with different safe levels, and filter-based detection [11]. For instance, the SD-v1.4 produces totally 6,403 exposed body parts among all resulting images on the NSFW-56k dataset, and our method reduces this number to 58 (NRR=99.1%).
6.1.1 NSFW Content Reduction
@@
We compare SafeGen and baselines in mitigating the generation of NSFW content across different adversarial prompts. The NRR quantifies the reduction of exposed body parts within images generated by SafeGen and baselines in comparison to the original SD model, where the exposure is determined by the NudeNet [44] classifier.
@@
Overall effectiveness.    Table I shows that SafeGen outperforms the baselines by achieving the highest average NRR of 95.6% across across all adversarial prompts. The baselines exhibit a range of NRR values, with the lowest being 8.5% (SLD (Weak)) to 72.1% (ESD), averaging at 49.8%. In addition, a visual comparison between SafeGen and these baselines provided by Figure 11 further demonstrates the effectiveness of SafeGen.
@@
We have three observations. First, external methods, on average, successfully remove 60.2% of NSFW content. However, due to the limitations of filters used in training data censorship or inference-stage filtering, particularly those involving less obvious content, out-of-distribution explicit content, and perturbations such as SneakyPrompt-P with pseudo words, may evade detection. Second, text-dependent mitigation can remove 70.7% NSFW content on average if we only consider those methods with the highest safety level, i.e., ESD and SLD (Max). While ESD manipulates the model weights to erase predefined text NSFW concepts, it may not account for all variations of such content or new content that evolve over time (e.g., porn stars’ names), leading to less effectiveness in NSFW removal. The difficult-to-enumerate challenge also limits the performance of SLD. Third, it is worthwhile to mention that SafeGen archives an impressive performance on the NSFW-56k dataset, i.e., 99.1% NRR. In contrast, the other baselines show different degrees of effectiveness, e.g., from 5.6% (SLD (Weak)) to 70.0% (safety filter). These outcomes suggest that the NSFW-56K dataset may serve as a challenging benchmark for future works in this domain.

### v2

@@ 4.4. System Integration
From a systematic view, the self-attention layer regulation method of SafeGen can seamlessly integrate other defenses as a complement. Our design boosts the compliance of unconditionally vision-only (i.e., text-agnostic) process within the classifier-free guidance (as illustrated in Equation (2)) without interfering with the conditionally text-dependent process. Hence, our method can collaborate with internal text-dependent countermeasures, particularly the guidance-based SLD (schramowski2023safe, ) to provide stronger protection that ensures safety for both conditional $\epsilon_{\mathtt{U^{*}}(z_{t},c,t)}$ and unconditional $\epsilon_{\mathtt{U^{*}}(z_{t},t)}$ denoising diffusion processes. Similarly, our method aligns well with ESD (gandikota2023erasing, ). Our evaluation of the complementary perspective is elaborated in §6.3.
@@
Figure 9. SafeGen effectively mitigates sexually explicit content yet retains the high-fidelity benign creation.
5. Experiment Setup
@@
We implement SafeGen using Python 3.8 and Pytorch 1.12 on a Ubuntu 22.04 server. All experiments are performed using an A100-40GB GPU (NVIDIA). SafeGen merely edits the self-attention layers of the U-Net module in Stable Diffusion models and can integrate with text-dependent methods. We follow previous work (yang2023sneakyprompt, ; gandikota2023erasing, ) to use Stable Diffusion (version 1.4) unless specified, more details including hyperparameters can be found in Appendix B.
5.1. Baselines
@@
We compare SafeGen with eight baselines, each exemplifying the latest anti-NSFW countermeasures. According to our taxonomy, these baselines can be divided into three groups: (1) N/A: where the original SD serves as the control group without any protective measures. (2) External Mitigation: involving safety filters to block inadvertently generated NSFW images (huggingface_safety_checker, ), although susceptible to bypassing by adversarial prompts; alternatively, conducting the training data censorship to minimize exposure to NSFW content and retraining the SD model using the censored data (SD-2-1, ), requiring substantial computation resources. (3) Internal Mitigation: involving representative text-dependent methods that steer the denoising diffusion process away from NSFW areas. Existing work either adopts guidance-based (schramowski2023safe, ) or model weights modification-based (gandikota2023erasing, ), but both are text-dependent and need predefined NSFW concepts. The details of these baselines are listed as follows:
  • $\bullet$
@@
    [N/A] SD: Stable Diffusion (SD-1-4, ), we follow previous work (yang2023sneakyprompt, ; gandikota2023erasing, ) to use the officially provided Stable Diffusion V1.4 (SD-1-4, ).
  • $\bullet$
@@
    [External Filter] SD with safety filter: we use the officially released image-based safety checker (huggingface_safety_checker, ) to examine its performance in detecting unsafe images.
  • $\bullet$
@@
    [External Censorship] SD-V2.1: Stable Diffusion V2.1, we use the official version (SD-2-1, ), which is retrained on a large-scale dataset censored by external filters.
  • $\bullet$
@@
    [Internal Text-Dependent] SLD: Safe Latent Diffusion, we adopt the officially pre-trained model (huggingface_SLD, ); our configuration examines its four safety levels, i.e., weak, medium, strong, and max.
  • $\bullet$
@@
    [Internal Text-Dependent] ESD: Erased Stable Diffusion, we follow its instruction (gandikota2023erasing, ), which erases the concept “nudity” and trains the model for 1000 epochs with learning rate 1e-5.
5.2. Evaluating Metrics
@@
We evaluate a T2I model’s ability in safe generation from two perspectives: (1) sexual explicitness mitigation, which is used to evaluate the model’s effectiveness in reducing sexually explicit content generation; and (2) benign content preservation, which is used to evaluate the model’s ability to preserve the high quality benign content generation. We use the following four metrics.
  • $\bullet$
@@
    [Sexually Explicit Mitigation] NRR [^3] Please note that SafeGen aims to suppress the generation of “sexually explicit” images, while an exact definition of “sexual explicitness” is difficult due to various sociological factors. We follow existing works (gandikota2023erasing, ; schramowski2023safe, ; SEGA, ; pham2023circumventing, ) to use “nudity” as a commonly-used quantifiable metric that detects “sexual explicitness”. We provide further discussion in §8.: We follow ESD (gandikota2023erasing, ), which uses nudity removal rate (NRR) [^3] Please note that SafeGen aims to suppress the generation of “sexually explicit” images, while an exact definition of “sexual explicitness” is difficult due to various sociological factors. We follow existing works (gandikota2023erasing, ; schramowski2023safe, ; SEGA, ; pham2023circumventing, ) to use “nudity” as a commonly-used quantifiable metric that detects “sexual explicitness”. We provide further discussion in §8. as a metric for assessing a T2I model’s efficacy in moderating sexually explicit content from images compared with the [N/A] original SD-V1.4 without safety mechanisms. NRR is calculated by NudeNet (nudenet, ). For each generated image, NudeNet first identifies exposed body parts like breasts or genitalia, it then aggregates the number of all identified parts as the total number of nude parts found in the image. The NRR refers to the difference in the number of detected nude parts between SafeGen or baseline methods and the SD-V1.4 model, a higher NRR indicates more effectiveness, meaning that more identified nude parts generated by the SD-V1.4 model have been successfully moderated. We first illustrate the overall effectiveness of SafeGen on different datasets by showing the NRR on total identified parts, we then show that SafeGen continuously outperforms baselines with a higher NRR on different nude parts.
  • $\bullet$
@@
    [Sexually Explicit Mitigation / Benign Preservation] CLIP Score: CLIP enables machines to interpret the relationships between images and their associated captions. Based on its significant zero-shot transferability, for each prompt, CLIP score computes the average cosine similarity between the given CLIP text embedding and its generated CLIP image embedding. In terms of benign generation, a higher score denotes that the T2I model can faithfully reflect the user’s prompt by way of images. In contrast, when confronted with a sexually explicit prompt, a lower score indicates the tested T2I model is safer as its generation deviates from the adversary’s desire.
  • $\bullet$
@@
    [Benign Preservation] LPIPS Score: The Learned Perceptual Image Patch Similarity (LPIPS) score (zhang2018perceptual, ) is another metric for evaluating the fidelity of generated images. LPIPS works by mimicking human visual perception, it captures the difference between detailed image features, such as texture and color. A lower score on the LPIPS score indicates that the two images are more visually similar.
  • $\bullet$
@@
    [Benign Preservation] FID Score: Different from the LPIPS focuses on the detailed comparison between two images, the Frechet Inception Distance (FID) score (parmar2022aliased, ) is a metric to compare the quality and fidelity between a set of created images and the other set of reference images. We evaluate the benign generated images’ quality of T2I models based on FID scores. A lower score on the FID score means that the two image sets’ distributions are more similar.
5.3. Adversarial and Benign Prompt Benchmark
@@
Our methodology is evaluated using a comprehensive benchmark that encompasses four different prompt datasets. To assess the effectiveness of SafeGen in reducing sexually explicit content generation, we utilize three adversarial prompt datasets, including the widely tested I2P dataset, along with our constructed SneakyPrompt and NSFW-56k datasets. Additionally, we employ a benign prompt dataset, COCO-2017, to evaluate SafeGen’s ability in maintaining high-fidelity benign generation.
  • $\bullet$
@@
    I2P: Inappropriate Image Prompts (I2P, ) consist of manually-tailored NSFW text prompts on lexica.art, from which we select all sex-related prompts, resulting in a total of 931 samples.
  • $\bullet$
@@
    SneakyPrompt: To evaluate the effectiveness of SafeGen against adaptive adversaries capable of generating sexually connotated prompts via optimization, we reproduce SneakyPrompt (yang2023sneakyprompt, ) and provide two versions of re-use prompt: i.e., SneakyPrompt-N with natural words, and SneakyPrompt-P with pseudo words.
  • $\bullet$
@@
    NSFW-56k: This dataset consists of 56k textual prompts that reflects real-world instances of sexual exposure (NSFW_image, ). We follow the CLIP Interrogator (Clip_Interrogator, ) to use BLIP2 (li2023blip2, ) to get multiple candidate text captions of a given pornographic image, then choose the best prompt with the highest CLIP score (CLIP, ) between image and text captions.
  • $\bullet$
@@
    COCO-25k: We follow prior works (qu2023unsafe, ; schramowski2023safe, ; gandikota2023erasing, ) to use MS COCO datasets prompts (from 2017 validation subset) for benign generation assessment. Each image within this dataset has been captioned by five human annotators, and the associated images were utilized as reference to gauge image fidelity.
6. Evaluation: Objective Metrics
@@
  • $\bullet$
    [RQ1] How effective is SafeGen in mitigating the sexually explicit generation from different types of adversarial prompts?
  • $\bullet$
@@
6.1. RQ1: Sexually Explicit Mitigation
We compare SafeGen with eight baselines, i.e., SD with different countermeasures, and show SafeGen outperforms all baselines in mitigating sexually explicit generation across two key metrics. First, we use the nudity removing rate (NRR) to show that SafeGen is effective in removing the explicit content, e.g., explicit body parts, among different adversarial prompts. Second, we use the CLIP score to show that SafeGen can reduce the text-to-image alignment between various adversarial prompts and their generation.
@@ 6.1. RQ1: Sexually Explicit Mitigation
Table 1. [RQ1-NRR] Performance of SafeGen on nudity removal rate compared with baselines on different adversarial prompt datasets.
Mitigation Method NRR (Nudity Removal Rate) $\uparrow$
@@
N/A Original SD 0% 0% 0% 0%
Censorship & Filter (External) SD-V2.1 64.9% 54.1% 47.5% 66.4%
Safety Filter 71.2% 71.4% 74.7% 72.9%
Text-dependent (Internal) ESD 84.2% 85.3% 63.9% 74.4%
SLD (Max) 81.8% 80.3% 82.6% 73.6%
SLD (Strong) 58.8% 55.8% 71.1% 50.5%
SLD (Medium) 30.6% 26.9% 44.7% 25.9%
SLD (Weak) 14.1% 5.2% 12.1% 8.5%
Text-agnostic SafeGen (Ours) 98.2% 98.0% 92.7% 99.4%
@@
Figure 10. [RQ1-NRR] We show the nudity removal rate (NRR) in the generated images classified as nudity by NudeNet (nudenet, ) compared to that from the original SD-V1.4 model. Our approach effectively reduces the explicit nudity content and outperforms all prior methods, i.e., SD-V2.1 (SD-2-1, ), ESD (gandikota2023erasing, ), SLD (schramowski2023safe, ) with different safety levels, and filter-based detection (huggingface_safety_checker, ). For instance, the SD-v1.4 produces totally 4,533 exposed body parts among all resulting images on the NSFW-56k dataset, and our method reduces this number to 27 (NRR=99.4%).
6.1.1. Nudity Content Reduction
@@
We compare SafeGen and baselines in mitigating the generation of sexually explicit content across different adversarial prompts. In line with ESD (gandikota2023erasing, ), we employ NRR to quantifies the reduction of exposed body parts within images generated by SafeGen and baselines in comparison to the original SD model, where the exposure is determined by the NudeNet (nudenet, ).
@@
Overall effectiveness. Table 1 shows that SafeGen outperforms the baselines by achieving the highest average NRR of 95.6% across across all adversarial prompts. The baselines exhibit a range of NRR values, with the lowest being 8.5% (SLD (Weak)) to 72.1% (ESD), averaging at 49.8%. In addition, a visual comparison between SafeGen and these baselines provided by Figure 15 further demonstrates the effectiveness of SafeGen.
@@
We have three observations. First, external methods, on average, successfully remove 60.2% of nude content. However, due to the limitations of filters used in training data censorship or inference-stage filtering, particularly those involving less obvious content, out-of-distribution explicit content, and perturbations such as SneakyPrompt-P with pseudo words, may evade detection. Second, text-dependent mitigation can remove 70.7% nude content on average if we only consider those methods with the highest safety level, i.e., ESD and SLD (Max). While ESD manipulates the model weights to erase predefined textual unsafe concepts, it may not account for all variations of such content or new content that evolve over time (e.g., porn stars’ names), leading to less effectiveness in nudity removal. The difficult-to-enumerate challenge also limits the performance of SLD. Third, it is worthwhile to mention that SafeGen archives an impressive performance on the NSFW-56k dataset, i.e., 99.1% NRR. In contrast, the other baselines show different degrees of effectiveness, e.g., from 5.6% (SLD (Weak)) to 70.0% (safety filter). These outcomes suggest that the NSFW-56K dataset may serve as a challenging benchmark for future works in this domain.
Different nude body parts. Figure 10 shows the results of SafeGen and baseline methods in reducing the generation of various exposed body parts, e.g., M-Breasts or F-Breasts, on the NSFW-56k dataset, where ‘M’ stands for male and ‘F’ stands for female. SafeGen achieves a 99.1% NRR for total exposed body parts, while the others are less effective on some body parts. For example, SLD (Strong) exhibits a 22.2% NRR for buttocks, and SD-V2.1 has a 26.0% NRR for belly, which indicates their limitation on undefined or unseen NSFW concepts. Moreover, a -16.9% NRR on buttocks caused by SLD (Weak) suggests some safe measures can unintentionally steer the denoising diffusion process towards unsafe regions. Due to the page limitation, we display the removal results on other three datasets in Appendix E (Figure 16).
6.1.2. Explicit Text-to-image Alignment Reduction
Table 2 shows the results of SafeGen and baselines in reducing the text-to-image alignment among different adversarial prompts, rendering findings from two perspectives:
[t]